40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  DRMI_1f_26April2017.pngIMG_7373.JPGIMG_7372.JPG
    Reply  Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking  20170428_DRMI.pdf
       Reply  Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking  
          Reply  Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking  LSC_sensingNoise.pdfIMG_7378.JPG
             Reply  Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking  
                Reply  Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking  
                   Reply  Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB DRMI_noArms_April30.pdfMICH_OLTF.pdfC1NB_disp_40m_MICH_NB_30_April_2017.pdf40m_REFL_RFPDs_efficiency.pdf
                      Reply  Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions 
                         Reply  Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions C1NB_disp_40m_MICH_NB_30_April_2017.pdf
                            Reply  Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions 
                               Reply  Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling OL_BS_ITM_comp.pdfC1NB_disp_40m_MICH_NB_8_May_2017.pdfC1NB_disp_40m_MICH_NB_10_May_2017.pdf
                                  Reply  Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling 
                                  Reply  Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise actuation.jpg
                                     Reply  Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                        Reply  Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                           Reply  Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey D010001-B_40m.pdfD000183-C8_40m.pdf
                                              Reply  Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling SRM_bypass_plus_CoilDriver.pdfliso.zip
                                                 Reply  Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                    Reply  Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                       Reply  Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate DAC_noise_model.pdfC1NB_disp_40m_MICH_NB_22_May_2017.pdf
                                                          Reply  Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises ITMX_deWhite_ch1_noise.pdf
                                                             Reply  Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises  coilDriverNoises.pdf
                                                                Reply  Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled deWhitening_consolidated.pdfdeWhitening_measurements.zipETMY_OL.png
                                                                   Reply  Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled OL_noiseInjection.pdf
                                                                      Reply  Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled 
       Reply  Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF PRC_angularFF.pdfPRC_TFs.pdf
Message ID: 12954     Entry time: Fri Apr 28 02:04:36 2017     In reply to: 12951     Reply to this: 12955   12957
Author: gautam 
Type: Update 
Category: General 
Subject: DRMI locking  

I got a couple of ~30min long DRMI lock stretches today. The settings I used are essentially the same as what I had back in November. Though we have since made some changes to the IMC RF signal chain, I guess it is not unreasonable that the LSC Demod phases that worked then work now as well. 

In the lock stretches, I did the following:

  • Took loop measurements for MICH, PRCL, SRCL
  • Turned on the sensing oscillator lines for error signal calibration
  • Tried turning on the analog whitening on AS55, REFL11 and REFL55. The latter two worked fine, but everytime I turned the REFL55 whitening on, I broke the lock. I'm also unable to acquire lock if I leave the whitening turned on all the time. The ADC overflow indicators also indicate frequent overflows when I turn the whitening on. Oddly, this seems to happen even if I turn the analog whitening gain to 0dB - the signals look well within the ADC range on dataviewer and DTT timeseries mode. Not sure what's going on here, I will investigate further tomorrow.
  • We should have some stretches where we can look at the possibility of seismic feedforward for some DRMI length DOFs.

On the side, I'm also looking at whether the PRC angular feedforward filters, last trained in October 2016, remain valid. Even post midnight, I am unable to lock the DRMI without turning on the FF, and looking at the POP QPD PIT and YAW signal spectra with the FF on vs FF off, there is definitely some improvement in the 1-4Hz band (plot to follow), question is whether we can do better and hence improve the DRMI duty cycle/ make the lock acquisition easier. To this end, I centered the beam on the POP QPD after locking and dither aligning the PRC on carrier, and have taken some data to look at.

So, much data analysis to follow - the idea is to put together a DRMI noise budget with Evan's NB code. For now, here are the uncalibrated control signal spectra.

Attachment 1: 20170428_DRMI.pdf  52 kB  | Hide | Hide all
20170428_DRMI.pdf
ELOG V3.1.3-